### v2

# Exploring the dynamical interplay between mass-energy equivalence, interactions and entanglement in an optical lattice clock

Abstract
We propose protocols that probe manifestations of the mass-energy equivalence in an optical lattice clock (OLC) interrogated with spin coherent and entangled quantum states. To tune and uniquely distinguish the mass-energy equivalence effects (gravitational redshift and second order Doppler shift) in such a setting, we devise a dressing protocol using an additional nuclear spin state. We then analyze the dynamical interplay between photon-mediated interactions and gravitational redshift and show that such interplay can lead to entanglement generation and frequency synchronization dynamics. In the regime where all atomic spins synchronize, we show the synchronization time depends on the initial entanglement of the state and can be used as a proxy of its metrological gain compared to a classical state. Our work opens new possibilities for exploring the effects of general relativity on quantum coherence and entanglement in OLC experiments.
Introduction.—Understanding the interplay between quantum mechanics (QM) and general relativity (GR) is a fundamental quest for modern science. Nevertheless to date, measurements capable of genuinely witnessing this simultaneous interplay have not been realized in tabletop experiments. A push forward towards this milestone is becoming feasible thanks to recent improvements in precision and accuracy of atomic clocks and interferometers. For example, matter-wave interferometers have been used for stringent tests on the equivalence principle [1, 2, 3, 4]. In parallel, the resolution of the gravitational redshift using spatially separated clocks has improved from meter scales [5, 6] to sub millimeters [7, 8]. Furthermore, coherence times for both clock and matter-wave interferometer are now sufficiently long [7, 9] to consider integrating clock and interferometer on a single platform. These developments open up unique opportunities to search for new physics [10, 1] that could help reconcile the seemingly contradictory predictions of QM and GR.
Figure 1: (a) Schematic of an optical lattice clock (OLC) embedded in the curved spacetime (metric $g_{\mu\nu}$) formed by the earth’s gravity. Mass-energy equivalence is the leading order GR correction that translates internal energy difference $\hbar\omega_{0}$ between $|e\rangle$ and $|g\rangle$ states into a difference in the rest mass of an atom $\hbar\omega_{0}/c^{2}$. Such type of correction generates second-order Doppler shift $E_{\rm SDS}$ and gravitational redshift $E_{\mathrm{GRS},j}$ to the clock transition (see Eq. (3) and Eq. (4)). (b) Schematic of probing the dynamical interplay between gravitational redshift and collective cavity-mediated interactions (see Eq. (5)).
Experimental developments have in parallel driven a great deal of theoretical effort towards the understanding of quantum dynamics with GR corrections. These progresses encompass analyses of relativistic corrections to Hamiltonians considered specifically in the context of neutral-atom and trapped-ion systems [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], tests of mass-energy equivalence with atoms in internal superposition states including predictions of energy-dependent phase shifts, loss of coherence and spin-motion coupling induced by gravitational time dilation [23, 24, 25, 26, 27, 28, 29, 30, 31], among others [32, 33, 34, 35, 36, 37, 38]. However, understanding the direct consequence of GR effects in more complex scenarios such as many-body systems, where particles can interact over the entire array, remains an outstanding problem.
In this work, we provide a first step in this direction by proposing near-term protocols to explore manifestations of single-atom GR effects, including the gravitational redshift (GRS) and the second-order Doppler shift (SDS), in the quantum many-body dynamics of an atomic ensemble. We take advantage of the state-of-the-art Wannier-Stark OLCs that interrogate large arrays of interacting particles under gravity [7, 39]. We analyze the GRS and the SDS specifically in Wannier-Stark OLCs, both governed by the same mechanism known as mass-energy equivalence. To distinctly characterize these GR effects and overcome the limitation that the gravitational redshift acts just as an effective magnetic field gradient in current OLCs, we firstly devise a dressing protocol using an additional nuclear spin state to tune and uniquely distinguish the mass-energy equivalence. Secondly, we propose a spectroscopic protocol to probe the dynamical modification of the GRS due to photon-mediated interactions in the array. Depending on the relative strength of the GRS to the photon-mediated interactions, the atomic phases dynamically evolve from their individual values, dictated by the local GRS, into a semi-local or global synchronized regime. In both cases we observe entanglement growth from an initial product state due to the interplay of the GRS and interactions, in contrast to the entanglement generated directly by gravitational interactions between massive objects [32, 33, 36, 37, 38]. Lastly, we analyze the interplay of the GRS and interactions when they act on an initially entangled state. This is achieved by studying the dependence of the global synchronization time on the entanglement content of the initial state. Remarkably we find the synchronization time can be used as a probe of the state’s metrological utility.
Mass-energy equivalence in OLCs.—We consider a single atom in earth’s gravity described by a curved spacetime metric $g_{\mu\nu}$ (see Fig. 1(a)). We perform an expansion of $g_{\mu\nu}$ in power series of $\phi/c^{2}$ [40, 41], with $\phi(Z)\approx g_{\rm LO}Z$ the Newtonian gravitational potential near the earth’s surface and $g_{\rm LO}$ the local gravitational acceleration. Following the treatment in Ref. [11, 14, 16, 18], one can obtain a single-atom Hamiltonian $\hat{H}_{A}$ accounting for the leading relativistic corrections:
$\hat{H}_{A}=\hat{H}_{\mathrm{point}}\bigg{(}M+\frac{\hat{H}_{I}}{c^{2}}\bigg{)}+O(c^{-4}).$ (1)
Here, $\hat{H}_{\mathrm{point}}(M)=Mc^{2}+\hat{H}_{0}(M)+\hat{H}_{\mathrm{other}}$ is the Hamiltonian of a point particle with mass $M$, $\hat{H}_{0}(M)=\hat{\mathbf{P}}^{2}/(2M)+M\phi$ contains the non-relativistic terms, and $\hat{H}_{\mathrm{other}}$ contains other GR corrections in center-of-mass coordinates negligible in our case (see  [^1] See Supplemental Material at [URL will be inserted by publisher] for details of GR corrections in an OLC, additional numerical simulations of superexchange interactions, as well as analytical derivation of synchronization time, including Ref. [40, 16, 18, 22, 14, 56, 57, 58, 59, 60, 44]). The key idea of Eq. (1) can be understood as the mass-energy equivalence, summarized by the replacement $M\to M+\hat{H}_{I}/c^{2}$ in $\hat{H}_{\mathrm{point}}$. OLCs feature an ultranarrow optical transition (clock transition) between two long-lived electronic states (excited state $|e\rangle$, ground state $|g\rangle$), which is described by the internal Hamiltonian $\hat{H}_{I}=\hbar\omega_{0}|e\rangle\langle e|$, with $\omega_{0}$ the clock transition frequency measured at the lab position $Z=0$ (see Fig. 1(a)). Since in an OLC $\hat{H}_{I}$ contains the largest observable energy scale compared to other terms, the mass-energy equivalence is the leading order GR correction. It translates into a difference in the rest mass of an atom in states $|e\rangle$ and $|g\rangle$: $M_{g}=M$, $\Delta M_{0}=M_{e}-M_{g}=\hbar\omega_{0}/c^{2}$. Note that in general the mass defect $\Delta M=\langle\hat{H}_{I}\rangle/c^{2}$ is not simply a fixed number, and its tunability (see Fig. 2) is an important tool to determine the relativistic origin of the mass defect.
We assume that in OLCs atoms are trapped in a magic-wavelength 1D lattice along the gravitational potential ($Z$ axis) [7, 39], where $|e\rangle$ and $|g\rangle$ states experience the same potential, $V(Z)=V_{Z}\sin^{2}(k_{L}Z)+M\phi$. Here $V_{Z}$ is the lattice depth, $k_{L}$ is the wave number of the lattice that sets the atomic recoil energy $E_{R}=\hbar^{2}k_{L}^{2}/2M$ and lattice spacing $a_{L}=\pi/k_{L}$. GR corrections to the optical lattice potential are negligible in our case (see [41]). In a tilted 1D lattice described by $V(Z)$, the motional eigenstates in the ground band are the so-called Wannier-Stark (WS) states $|W_{j}\rangle$, with $j$ the $Z$-lattice site index where the WS state is centered at [39]. Assuming the radial degrees of freedom are also confined to the lowest ground state by an additional 2D lattice, with lattice depths $V_{X,Y}$, the eigenenergies of WS states are given by $E_{j}=Mg_{\rm LO}a_{L}j+E_{\mathrm{band}}$, where $E_{\mathrm{band}}\approx\sum_{\alpha=X,Y,Z}E_{R}(\sqrt{V_{\alpha}/E_{R}}-1/4)$ [42] is the ground band zero-point energy. The GR corrections due to mass-energy equivalence is given by
$\hat{H}_{\mathrm{corr}}=\sum_{j}(E_{\mathrm{GRS},j}+E_{\mathrm{SDS}})|e,W_{j}\rangle\langle e,W_{j}|,$ (2)
with $E_{\mathrm{GRS},j}$ the gravitational redshift (GRS) and $E_{\mathrm{SDS}}$ the second-order Doppler shift (SDS). Their orders of magnitude are discussed below for 87Sr atoms.
Applying the mass-energy equivalence to the gravitational potential energy $M\phi$, we get the GRS (or gravitational time dilation) between $|e\rangle$ and $|g\rangle$ states ($\Delta M=\Delta M_{0}$),
$E_{\mathrm{GRS},j}=\frac{\Delta M}{M}\langle W_{j}|M\phi|W_{j}\rangle=\hbar\omega_{0}\frac{g_{\rm LO}a_{L}j}{c^{2}}.$ (3)
The GRS leads to a gradient of frequency shifts across the lattice. For example, the fractional frequency difference for nearest-neighbor lattice sites is just $4.4\times 10^{-23}$, while it is at the order of $10^{-19}$ for $1$ mm spatial separation as recently observed [7, 8].
The contribution of mass-energy equivalence in the kinetic energy leads to a local modification of the zero-point energy known as SDS (or motional time dilation in special relativity) between $|e\rangle$ and $|g\rangle$ states,
$E_{\mathrm{SDS}}=-\frac{\Delta M}{M}\frac{\langle W_{j}|\hat{\mathbf{P}}^{2}|W_{j}\rangle}{2M}=-\frac{\hbar\omega_{0}}{2Mc^{2}}E_{\mathrm{band}}.$ (4)
The magnitude of $E_{\mathrm{SDS}}$ increases with the lattice depth. For example, a deep lattice with $V_{X,Y,Z}=300E_{R}$ leads to fractional frequency shift $-4.5\times 10^{-21}$. Corrections in the kinetic energy can also lead to a modification of the WS wave functions for $|e\rangle$ atoms, while they play a negligible role compared to $E_{\mathrm{GRS},j}$ and $E_{\rm SDS}$.
Figure 2: Tuning mass-energy equivalence via dressed states. (a) Schematic of dressing the clock transition with another nuclear spin. The left panel show the application of the dressing laser, and the right panel show the new clock transition in the dressed basis. (b) The tunability of the mass defect $\Delta M$ and gravitational redshift $\omega_{\mathrm{GRS}}$ as a function of dressing parameter $\delta/\Omega$. $\Delta M_{0}$ and $\omega_{\mathrm{GRS},0}$ are the corresponding values without dressing. (c) The tunability of cavity-mediated interactions (see $\hat{H}_{\rm cGR}$ in Eq. (5)) as a function of dressing parameters $\delta/\Omega$ and nuclear spin level $m_{F}$. Heisenberg interaction ($J_{\perp}=J_{z}$) can be reached with $m_{F}=3/2$.
Tuning and distinguishing GR effects.—In standard OLCs, the effects of GRS might be mimicked by a weak magnetic field gradient. To provide further evidence of genuine GR effects beyond ruling out all possible systematics, one approach is to simultaneously observe $E_{\mathrm{GRS},j}$ and $E_{\rm SDS}$ in the same system. This could be achieved in next-term OLCs by populating higher motional bands if the systematic uncertainty of lattice Stark shifts [43] is suppressed below $10^{-20}$.
[…]
This protocol is feasible thanks to the fact that the clock transitions between different nuclear spins are frequency resolved due to magnetic Zeeman shifts. We also assume all other dynamical frequencies are smaller than the dressed state energy gap $\sqrt{\Omega^{2}+\delta^{2}}$ and the Zeeman shifts between nuclear spins. To guarantee the matching of laser phases for each atom, the dressing beam and the clock beam should be co-propagating. Moreover, spatial inhomogeneities in atomic detunings $\delta$ and in the dressing laser Rabi frequency $\Omega$ might obscure the effects of gravitational redshift. For a mHz gravitational redshift arising from a cm-scale spatial separation, the spatial variations of $\delta(Z)$, $\Omega(Z)$ and other source of perturbations need to be suppressed below $10^{-4}$ Hz. The requirement for $\delta(Z)$ is attainable based on the parameters in Ref. [7]. The requirement for $\Omega(Z)$ could be achievable using a cavity to stabilize the spatial mode of the dressing laser. One can also circumvent the stringent requirement for $\Omega(Z)$ by averaging the transition frequency of $|g,m_{F}\rangle\leftrightarrow|-\rangle$ and $|g,m_{F}\rangle\leftrightarrow|+\rangle$ (see Appendix), while sacrificing the tunability of mass defect ($\Delta M=\Delta M_{0}/2$).
Many-body dynamics.— After providing a recipe to distinguish genuine GR effects in OLCs, we further explore their manifestations in quantum many-body dynamics. We consider photon-mediated interactions generated by placing the atoms in an optical cavity [44, 45], in a regime where atomic contact interactions are controlled to be much weaker. The interplay between photon-mediated interactions and the GRS is described by the following Hamiltonian (see Fig. 1(b) and Appendix):
$\hat{H}_{\rm cGR}/\hbar=J_{\perp}\hat{\bf S}\cdot\hat{\bf S}+(J_{z}-J_{\perp})\hat{S}^{z}\hat{S}^{z}+\omega_{\mathrm{GRS}}\sum_{j}j\hat{\mathcal{S}}^{z}_{j},$ (5)
where $\hbar\omega_{\mathrm{GRS}}=(\Delta M)g_{\rm LO}a_{L}$ is the GRS between nearest neighbor sites, $J_{\perp}$ and $J_{z}$ are the collective exchange and Ising couplings. Here $\hat{\mathcal{S}}_{j}^{x,y,z}$ are collective spin operators summed over all atoms at the same height $ja_{L}$, and $\hat{S}^{x,y,z}=\sum_{j}\hat{\mathcal{S}}_{j}^{x,y,z}$. Based on Eq. (5), a magnetic field gradient will in principle give rise to similar single-atom inhomogeneities in the Hamiltonian, but we can tell them apart using the dressing scheme. We drop the GR corrections for interaction terms since they are negligible in our case (see [41]). While the use of a single nuclear spin state restricts the cavity exchange interactions to a single polarization mode ($J_{z}=0$ only), the dressing to another nuclear spin allows for coupling two polarization modes of the cavity (see Appendix), which enhances the tunability of $\hat{H}_{\rm cGR}$ and realizes collective Heisenberg interactions ($J_{\perp}=J_{z}$) as shown in Fig. 2(c). In the following, we mainly focus on the case of $J_{\perp}=J_{z}$, since the $\hat{\bf S}\cdot\hat{\bf S}$ term becomes a constant and does not alter entanglement in the fully symmetric manifold. This requirement is unnecessary for observing frequency synchronization.
We propose to initialize all the atoms in the state $(|\uparrow\rangle+|\downarrow\rangle)/\sqrt{2}$ ($\pi/2$ pulse between the dressed clock levels), perform time evolution under Hamiltonian $\hat{H}_{\rm cGR}$ (Eq. (5)), and then measure the phase shift of $\langle\hat{\mathcal{S}}^{+}_{j}\rangle$ for every $Z$-lattice site, resulting in frequency shift $\omega_{j}(t)=\tan^{-1}\left[\langle\hat{\mathcal{S}}^{y}_{j}\rangle/\langle\hat{\mathcal{S}}^{x}_{j}\rangle\right]/t$ as a function of evolution time (see Fig. 3(a)). It can be observed by the application of a $\pi/2$ pulse followed by local population measurements. Without interactions or in the case of short interrogation times, one expects to observe the GRS, $\omega_{j}=j\omega_{\mathrm{GRS}}$, as reported in Ref. [7]. In the interaction dominated regime, the GRS persists only for a time scale shorter than the atomic interaction time scale. Beyond this period, the frequencies become synchronized due to interaction locking and reach $\omega_{j}\approx 0$ at synchronization time $t_{\mathrm{syn}}$ (see Fig. 3(b)). Without loss of generality, the numerical simulations in Fig. 3 and Fig. 4 are based on exact diagonalization for $N=16$ atoms with one atom per $Z$-lattice site.
[…]
Figure 4: Interplay between entanglement and GRS. (a) One-axis twisting (OAT) interactions $\hat{U}_{\mathrm{OAT}}(Q)=\exp(-iQ\hat{S}^{z}\hat{S}^{z}/N)$ and rotations $\hat{R}_{x}^{\theta}=\exp(-i\theta\hat{S}^{x})$ are first applied to generate a spin squeezed initial state (at $Q/2\pi=0.6$), followed by unitary evolution under $\hat{H}_{\rm cGR}$. (b) The synchronization time, $t_{\mathrm{syn}}$ depends on the orientation of the spin squeezed state determined by $\theta$. (c) $\Delta t_{\mathrm{syn}}$ (marked in (b)) as a function of OAT shearing strength $Q$. We show $C_{yz}=4\mathrm{max}_{\theta}\mathrm{Cov}(y,z)$ is approaching the quantum Fisher information $F_{Q}$ for spin squeezed states (inset). We compare the numerical simulations (blue points) for $N=16$ atoms under $\hat{H}_{\rm cGR}$ ($J_{\perp}=J_{z}$) with Eq. (6) (black dashed lines).
The emergent synchronization is the result of many-body gap protection also observed in prior experiments [46, 47, 44, 48, 49, 50], which arises when $\omega_{\rm split}\ll\Delta E$. Here $\omega_{\rm split}=(N_{s}-1)\omega_{\rm GRS}$ is the maximum redshift in the array, with $N_{s}$ the number of Z-lattice sites, and $\Delta E$ is the many-body gap due to Heisenberg couplings. On the contrary, in the regime $\omega_{\rm split}\sim\Delta E$, the gap cannot maintain global synchronization (see Fig. 3(c)). Using a spin wave analysis one obtains $\Delta E=NJ_{\perp}$ and $NJ_{\perp}t_{\mathrm{syn}}=\pi$ for $J_{\perp}=J_{z}$, where $N$ corresponds to the total atom number in the array [41]. For $N\sim 10^{5}$ 87Sr atoms, one can achieve $NJ_{\perp}/2\pi\sim$ Hz ($\omega_{\rm split}/NJ_{\perp}\sim 10^{-3}$ for cm-scale separation) [44], leading to a synchronization time scale ($\sim 1$s) within reach of current experiments.
Furthermore, we find that the simultaneous presence of single-atom GRS and photon-mediated interactions can lead to quantum entanglement as shown in Fig. 3(d,e). In fact, in the regime $\omega_{\rm split}\ll\Delta E$, the projection of the wave function into the fully symmetric manifold imposed by the many-body gap transforms the single-particle term into an effective one-axis twisting (OAT) [51, 52] interaction term $\chi\hat{S}^{z}\hat{S}^{z}$, with $\chi\sim\omega_{\rm split}^{2}/[N(\Delta E)]$ (see Ref. [53] where the splitting is generated by a different mechanism). In this case, entanglement builds up for $t>t_{\rm syn}$, as witnessed by a squeezing parameter [52], $\xi^{2}\equiv\min_{\varphi}N(\Delta S_{\varphi}^{\perp})^{2}/|\langle\hat{\mathbf{S}}\rangle|^{2}<1$ (see Fig. 3(d)). Here, $(\Delta S_{\varphi}^{\perp})^{2}$ is the variance of spin noise along an axis perpendicular to the collective spin $\langle\hat{\mathbf{S}}\rangle$. A faster growth of entanglement can be seen in the regime $\omega_{\rm split}\sim\Delta E$ (see Fig. 3(e)). Since the entanglement in this case is not captured by spin squeezing, instead we characterize the entanglement by the normalized Rényi entropy $\tilde{S}_{N/2}=-2\log_{2}(\mathrm{tr}(\hat{\rho}_{N/2}^{2}))/N$, where $\hat{\rho}_{N/2}$ is the reduced density matrix by taking partial trace over half of the system. The entanglement builds up at a time scale $t_{\rm split}\sim\pi/\omega_{\rm split}$ in this case, which might be due to population transfer to highly entangled states in manifolds of lower total spin. For the implementation of entanglement generation, $\omega_{\rm split}/NJ_{\perp}\sim 0.1-1$ is achievable for $10$ cm -- $1$ m separation and $NJ_{\perp}/2\pi\sim 0.1$ Hz. Therefore, the time scale for entanglement growth is at the order of $10$s and thus requires careful suppression of systematics.
To study the effects of the GRS on quantum entanglement, we consider the scenario with entangled initial states, such as the ones generated using cavity induced OAT interactions [54, 55], $\hat{U}_{\mathrm{OAT}}(Q)=\exp(-iQ\hat{S}^{z}\hat{S}^{z}/N)$, where $Q$ is the shearing strength (Fig. 4(a)). The squeezing direction corresponds to the direction with minimum value of $(\Delta S_{\varphi}^{\perp})^{2}$, which can be controlled by performing a rotation $\hat{R}_{x}^{\theta}=\exp(-i\theta\hat{S}^{x})$ as shown in Fig. 4. As demonstrated in Ref. [55], entangled state preparation time scales in an OLC are at the order of $10$ ms, so it is reasonable to ignore GRS in initial state preparation. We use $|\psi_{0}\rangle$ to denote the state after OAT interactions, and $|\psi(\theta)\rangle$ for the state after the $\hat{R}_{x}^{\theta}$ rotation.
In Fig. 4(b), we show that it is possible to control $t_{\mathrm{syn}}$ below or above the value of a product initial state $t_{\mathrm{syn},0}$ ($Q=0$, as obtained in Fig. 3) depending on the rotation $\hat{R}_{x}^{\theta}$. The ratio $t_{\mathrm{syn}}/t_{\mathrm{syn},0}$ under $\hat{H}_{\rm cGR}$ ($J_{\perp}=J_{z}$) can be understood using the following analytic result [41],
$\frac{t_{\mathrm{syn}}}{t_{\mathrm{syn},0}}=1-\frac{2}{\pi}\arctan\bigg{[}\frac{{\rm Cov}(y,z)}{(N-1)\langle\psi_{0}|\hat{S}^{x}|\psi_{0}\rangle}\bigg{]},$ (6)
where ${\rm Cov}(\alpha,\beta)\equiv\langle\psi(\theta)|(\hat{S}^{\alpha}\hat{S}^{\beta}+\hat{S}^{\beta}\hat{S}^{\alpha})|\psi(\theta)\rangle-2\langle\psi(\theta)|\hat{S}^{\alpha}|\psi(\theta)\rangle\langle\psi(\theta)|\hat{S}^{\beta}|\psi(\theta)\rangle$, with $\alpha,\beta=x,y,z$. The tunability of $t_{\mathrm{syn}}$ is due to the $\theta$-dependence of ${\rm Cov}(y,z)$. The tunable range $\Delta t_{\mathrm{syn}}\equiv\max_{\theta}t_{\mathrm{syn}}-\min_{\theta}t_{\mathrm{syn}}$ can be used as a measure of entanglement (see Fig. 4(c)), since $C_{yz}=4\max_{\theta}{\rm Cov}(y,z)$ is approaching the quantum Fisher information $F_{Q}$ [52], which corresponds to the maximal eigenvalue of the matrix $F_{Q,\alpha\beta}=2{\rm Cov}(\alpha,\beta)$.
Conclusion and outlook.—We discussed protocols accessible in OLCs that explore how the single-atom GR effects modify many-body dynamics generated by photon-mediated interactions. A similar interplay should be observable with atomic superexchange interactions. While so far we have mostly focused on highly localized atomic arrays, generalizations to the case of itinerant particles where motion and other GR corrections also become relevant, will open unique opportunities for testing the basic tenets of GR when extended into the complex quantum domain.
Acknowledgements.
We thank James Thompson, Raphael Kaubruegger, Dylan Young, Alexander Aeppli, Michael Werner and Torsten Zache for useful discussions. This work is supported by the Sloan Foundation, the Simons Foundation and the Heising-Simons Foundation, NSF JILA-PFC PHY-2317149 and NSF OMA-2016244 (QLCI) grants, and by NIST. VML and KH acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Project-ID 274200144 – SFB 1227 (project A09) and Project-ID 390837967 - EXC 2123.
Appendix A Appendix: Derivations of the dressing scheme
Construction of dressed states.—Here we explain the dressing scheme in Fig. 2 in more detail. Due to magnetic Zeeman shifts, the clock transitions between different nuclear spins are frequency resolved, thus we can restrict the dynamics within three internal levels $|e,m_{F}\rangle$, $|g,m_{F}\rangle$ and $|g,m_{F}-1\rangle$. A dressing laser (with Rabi frequency $\Omega$, laser frequency $\omega_{d}$) is used to couple $|e,m_{F}\rangle$ with $|g,m_{F}-1\rangle$ states (see Fig. 2(a)), leading to the following Hamiltonian,
$\hat{H}_{\rm dress}/\hbar$ $=\omega_{0}\hat{\mathcal{P}}_{e,m_{F}}-\omega_{Z}\hat{\mathcal{P}}_{g,m_{F}-1}$ (7)
$+\frac{\Omega}{2}(|e,m_{F}\rangle\langle g,m_{F}-1|e^{-i\omega_{d}t}+h.c.),$
where $\omega_{Z}$ is the Zeeman shift between $|g,m_{F}\rangle$ and $|g,m_{F}-1\rangle$ states, and we set the energy of $|g,m_{F}\rangle$ state to 00. The projection operators are defined by $\hat{\mathcal{P}}_{\psi}=|\psi\rangle\langle\psi|$. We then perform a unitary transformation $\hat{U}=\exp(-i\omega_{d}t\,\hat{\mathcal{P}}_{g,m_{F}-1})$ to rewrite the Hamiltonian in the rotating frame of the dressing laser,
$\hat{H}_{\rm dress}^{\prime}/\hbar$ $=\omega_{0}\hat{\mathcal{P}}_{e,m_{F}}+(\omega_{0}+\delta)\hat{\mathcal{P}}_{g,m_{F}-1}$ (8)
$+\frac{\Omega}{2}(|e,m_{F}\rangle\langle g,m_{F}-1|+h.c.),$
where $\delta=\omega_{d}-\omega_{0}-\omega_{Z}$ is the detuning of the dressing laser. The eigenstates of this Hamiltonian are given by $|+\rangle=C_{1}|e,m_{F}\rangle+C_{2}|g,m_{F}-1\rangle$, $|-\rangle=-C_{2}|e,m_{F}\rangle+C_{1}|g,m_{F}-1\rangle$ with energy $E_{\pm}/\hbar=\omega_{0}+\delta/2\pm\sqrt{\Omega^{2}+\delta^{2}}/2$. These are the dressed states defined in the main text with coefficients $C_{1}=(1-\delta/\sqrt{\Omega^{2}+\delta^{2}})^{1/2}/\sqrt{2}$ and $C_{2}=(1+\delta/\sqrt{\Omega^{2}+\delta^{2}})^{1/2}/\sqrt{2}$.
Tunability of GR effects.—Firstly we analyze the effects of the dressing scheme on the GR effects. We focus on the GRS described by the following Hamiltonian,
$\hat{H}_{\rm GRS}=(\Delta M_{0})\phi\hat{\mathcal{P}}_{e,m_{F}}=$ $\,(\Delta M_{0})\phi\bigg{(}|C_{2}|^{2}\hat{\mathcal{P}}_{-}+|C_{1}|^{2}\hat{\mathcal{P}}_{+}$ (9)
$+\Big{(}C_{1}C_{2}^{*}|+\rangle\langle-|\,+h.c.\Big{)}\bigg{)},$
where $\Delta M_{0}=\hbar\omega_{0}/c^{2}$ is the mass defect without the dressing protocol. As shown in Fig. 2(a), we probe the transition between $|g,m_{F}\rangle$ and |ket|$\rangle$| - ⟩. Consider the case that the energy gap $\sqrt{\Omega^{2}+\delta^{2}}$ between |+ket|$\rangle$| + ⟩ and |ket|$\rangle$| - ⟩ states due to $\hat{H}_{\rm dress}$ is much larger than $\omega_{0}\phi/c^{2}$, the coupling between |+ket|$\rangle$| + ⟩ and |ket|$\rangle$| - ⟩ states in $\hat{H}_{\rm GRS}$ are suppressed significantly. Therefore, to the leading order, when projected into the relevant $|g,m_{F}\rangle$ and |ket|$\rangle$| - ⟩ states we can write $\hat{H}_{\mathrm{GRS}}$ as:
$\hat{H}_{\mathrm{GRS}}\approx(\Delta M_{0})\phi|C_{2}|^{2}\hat{\mathcal{P}}_{-}=(\Delta M)\phi\hat{\mathcal{P}}_{-},$ (10)
which is equivalent to a modification of the mass defect (see Fig. 2(b)),
$\Delta M=|C_{2}|^{2}\Delta M_{0}.$ (11)
The same analysis applies to the SDS.
The tunability of gravitational redshift due to the dressing scheme allows us to distinguish GRS from a magnetic field gradient. In the dressing scheme, $\hat{H}_{\rm GRS}$ leads to a position-dependent correction of the dressed state energy $E_{-}$,
$E_{-}(Z)=E_{-}+\frac{\hbar\omega_{0}g_{\rm LO}}{c^{2}}|C_{2}|^{2}Z,$ (12)
which can be resolved via clock spectroscopy in the effective two level system formed by $|g,m_{F}\rangle$ and |ket|$\rangle$| - ⟩ states. While for a small magnetic field gradient term adding on top of a constant magnetic field, we have $\omega_{0}(Z)=\omega_{0}+(\eta_{e}-\eta_{g})m_{F}Z$, $\omega_{Z}(Z)=\omega_{Z}+\eta_{g}Z$, where $\eta_{e}=-\mathcal{G}_{{}^{3}P_{0}}\mu_{B}\partial_{Z}B$, $\eta_{g}=-\mathcal{G}_{{}^{1}S_{0}}\mu_{B}\partial_{Z}B$, with $\mathcal{G}_{{}^{3}P_{0}}$ and $\mathcal{G}_{{}^{1}S_{0}}$ representing the Landé g-factors, and $\mu_{B}$ is the Bohr magneton. So the position-dependent correction due to magnetic field gradients is given by
$E_{-}(Z)=E_{-}+\bigg{[}|C_{2}|^{2}(\eta_{e}-\eta_{g})m_{F}-|C_{1}|^{2}\eta_{g}\bigg{]}Z.$ (13)
Since $\eta_{g}\neq 0$, we find different dependence by varying $\delta$ compared to the GRS (see Eq. (12)). The reason is that different ground-state nuclear spins have the same mass but different Zeeman shifts.
Figure 5: Schematic of the cavity-mediated interactions in the dressed basis. The cavity axis is along $Z$ and the quantization axis is along magnetic field $B$ labeled in the plot. The additional nuclear spin state in the dressing protocol allows for coupling to two polarization modes in the cavity and enhancing the tunability of the cavity-mediated interactions.
Tunability of cavity-mediated interactions.—Secondly we analyze the effects of the dressing protocol on the cavity-mediated interactions. We focus on multilevel alkaline earth atoms with the quantization axis for nuclear spins perpendicular to the cavity axis (see Fig. 5). In this case, the two polarization modes supported by the cavity can drive the $\pi$ transition and the linear combination of $\sigma^{+}$ and $\sigma^{-}$ transitions, so we can define the multilevel raising operators for these two polarization modes, $\hat{\Pi}^{+}=\sum_{jm}C^{0}_{m}|e,m\rangle_{j}\langle g,m|$, $\hat{\Sigma}^{+}=\sum_{jm}\frac{i}{\sqrt{2}}(C^{-1}_{m}|e,m-1\rangle_{j}\langle g,m|+C^{+1}_{m}|e,m+1\rangle_{j}\langle g,m|)$, where $j$ is the label of atoms, $m$ is the label of nuclear spins, and $C^{\sigma}_{m}\equiv\langle F,m;1,\sigma|F,m+\sigma\rangle$ are the Clebsch-Gordan coefficients. The photon-mediated exchange interactions for multilevel alkaline earth atoms take the following form [61, 62],
$\hat{H}_{c}/\hbar=\chi(\hat{\Pi}^{+}\hat{\Pi}^{-}+\hat{\Sigma}^{+}\hat{\Sigma}^{-}),$ (14)
[…]
$\hat{H}_{c}/\hbar\approx J_{\perp}\hat{S}^{+}\hat{S}^{-}+J_{z}\bigg{(}\frac{N}{2}+\hat{S}^{z}\bigg{)}^{2},$ (15)
where
$J_{\perp}=\chi(C^{0}_{m_{F}})^{2}|C_{2}|^{2},\quad J_{z}=\chi\frac{(C^{+1}_{m_{F}-1})^{2}}{2}|C_{1}|^{2}|C_{2}|^{2}.$ (16)
The tunability of $J_{\perp}$ and $J_{z}$ is shown in Fig. 2(c). Eq. (5) describing the interplay between gravitational redshift and cavity-mediated interactions is a combination of Eq. (10) and Eq. (15) in the Wannier-Stark basis.
The physical meaning of the $J_{\perp}$ term is the spin exchange interactions between $|g,m_{F}\rangle$ and |ket|$\rangle$| - ⟩ state via the $\hat{\Pi}^{+}\hat{\Pi}^{-}$ process, i.e. an atom in |ket|$\rangle$| - ⟩ state emits a virtual photon into the cavity and flips to $|g,m_{F}\rangle$ state ($\hat{\Pi}^{-}$), and another atom in $|g,m_{F}\rangle$ state absorbs the same photon and flips to |ket|$\rangle$| - ⟩ state ($\hat{\Pi}^{+}$). The physical meaning of $J_{z}$ term is the collective frequency shift of |ket|$\rangle$| - ⟩ state via the $\hat{\Sigma}^{+}\hat{\Sigma}^{-}$ process, i.e. an atom in |ket|$\rangle$| - ⟩ state absorbs a photon from the dressing laser beam and emits it into the cavity while staying in the same state ($\hat{\Sigma}^{-}$), then another atom in |ket|$\rangle$| - ⟩ state absorbs the same cavity photon and emits it back to the dressing laser beam ($\hat{\Sigma}^{+}$).
Experimental considerations.—Finally we focus on the experimental requirements of the dressing scheme. For 87Sr atoms, we consider the Zeeman shifts between nuclear spin states to be at the order of $10^{2}$ Hz. In order to frequency resolve a single transition between nuclear spin states, we have $\Omega/2\pi,\delta/2\pi\sim 10$ Hz. Considering $\chi N/2\pi$ at the order of Hz for clock transition as shown in Ref. [44], and GRS at the order of mHz for $1$ cm spatial separation, the validity of the dressing scheme is ensured.
As for experimental implementation, spatial inhomogeneities exist for the atomic frequency and the parameters $\delta$ and $\Omega$ in $E_{-}$. The effects of GRS will be washed out if the inhomogeneities are much larger than the value of GRS. For $1$ cm spatial separation, the mHz scale of GRS requires to control the inhomogeneities below $10^{-4}$ Hz for direct observation.
For the atomic frequencies and detuning $\delta$, the leading order contributions are from spatial inhomogeneties in the magnetic field. One can suppress first order Zeeman shifts by probing opposite nuclear spin states and calculating the averaged frequency, which is attainable based on the parameters in Ref. [7]. This approach allows for cancellation of first order Zeeman shifts up to shot-to-shot fluctuations, and residue effects of the magnetic field can be distinguish from GRS based on the dressing scheme.
For the dressing laser Rabi frequency $\Omega$, the leading order contributions are from the spatial profile of the laser beam. If we denote the modification of $\Omega$ as $\Delta\Omega$, the change of $E_{\pm}$ is given by
$\Delta E_{\pm}/\hbar=\pm\frac{1}{2}\sqrt{\Omega^{2}+\delta^{2}}\bigg{(}1+\frac{(\Delta\Omega)\Omega}{\Omega^{2}+\delta^{2}}\bigg{)}.$ (17)
The suppression of inhomogeneties in $E_{-}$ requires $\Delta\Omega/2\pi<10^{-4}~{}$z, which is equivalent to $\Delta\Omega/\Omega<10^{-5}$. In principle, this requirement is achievable using an ultrastable cavity, which allows for precise control of the spatial mode of the dressing laser. Notice that $\Delta E_{+}+\Delta E_{-}=0$, an alternative approach to reduce this requirement is to average the transition frequency of $|g,m_{F}\rangle\leftrightarrow|-\rangle$ and $|g,m_{F}\rangle\leftrightarrow|+\rangle$,
$\frac{E_{-}(Z)+E_{+}(Z)}{2}=\hbar\omega_{0}+\frac{\hbar\delta}{2}+\frac{1}{2}\frac{\hbar\omega_{0}}{c^{2}}\phi(Z).$ (18)
In this way, the averaged transition frequency becomes independent of $\Omega$. Even though one sacrifices the full tunability of GRS, it is still changed to half of its value without dressing.